# Prot2Token: A Unified Framework for Protein Modeling via Next-Token Prediction

### v1

@@ Abstract
The diverse nature of protein prediction tasks has traditionally necessitated specialized models, hindering the development of broadly applicable and computationally efficient Protein Language Models (PLMs). In this work, we introduce Prot2Token, a unified framework that overcomes these challenges by converting a wide spectrum of protein-related predictions—from sequence-level properties and residue-specific attributes to complex inter-protein interactions—into a standardized next-token prediction format. At its core, Prot2Token employs an autoregressive decoder, conditioned on embeddings from pre-trained protein encoders and guided by learnable "task tokens", to perform diverse predictions. This architecture uniquely facilitates multi-task learning, enabling a single model to master numerous tasks with improved efficiency. We present extensive experimental validation across a variety of benchmarks, demonstrating Prot2Token’s strong predictive power in different types of protein-prediction tasks. Key results include significant speedups (e.g.,  1000x over AlphaFold2 with MSA) and performance often matching or exceeding specialized approaches. Beyond that, we introduce an auxiliary self-supervised decoder pre-training approach to improve spatially sensitive task performance. Prot2Token thus offers a significant step towards a versatile, high-throughput paradigm for protein modeling, promising to accelerate biological discovery and the development of novel therapeutics. The code is available at GitHub.
@@ 1 Introduction
While PLMs have significantly advanced protein-prediction tasks, current models require task-specific specialization after pre-training [5, 6]. This reliance on separate modules for distinct tasks leads to inefficient computational resource use and limited scalability. Most PLMs undergo post-training alignment with specialized predictor architectures for individual tasks, requiring independent training and fine-tuning—a time-consuming and resource-intensive approach [7]. A unified model capable of efficiently handling diverse protein-prediction tasks would overcome this limitation, streamlining protein function prediction and enhancing its accessibility for real-world applications.
To the best of our knowledge, despite the emergence of foundation models for proteins, no comprehensive framework exists to systematically align them across a broad spectrum of protein prediction tasks. Instead, researchers often modify existing foundation models to suit particular applications [8], such as predicting 3D protein structures from sequences using customized techniques [9, 10]. One key limitation is that most existing models are based on BERT-style architectures [11], while effective for providing meaningful representation, lack the flexibility needed for diverse and controllable prediction capabilities. In natural language processing (NLP), the transition from BERT-style models to autoregressive GPT-style models has enabled more dynamic and human-understandable instructions (prompts) to control the generation process and therefore, handling a diverse set of predictions within the NLP domain [12, 13]. A similar paradigm shift is necessary in protein research, moving beyond static encoders toward more advanced generative AI approaches that provide more comprehensive predictive capabilities.
@@
To address these limitations, this work takes a significant step toward unifying diverse protein-related prediction tasks within a single comprehensive framework (Figure 1). We introduce a universal protocol for tokenizing different protein-prediction tasks, enabling a general autoregressive transformer predictor to leverage existing BERT-style PLMs. This unified autoregressive predictor, guided by next-token prediction loss, demonstrates strong generality across multiple protein-prediction task categories, including protein-level, residue-level, and protein-protein interaction-level tasks. We illustrate its versatility through extensive evaluation on diverse examples such as kinase phosphorylation site prediction, protein-ligand binding site prediction, protein 3D structure prediction, and protein mutation stability assessment. Furthermore, our framework inherently supports multi-task learning, and we provide initial analyses demonstrating synergistic performance improvements when related tasks are trained jointly. For certain specialized tasks, such as predicting binding sites, we show that initializing the decoder through self-supervised pre-training significantly boosts performance. Specifically, for protein-ligand binding site prediction, we further analyzed the learned token representations, revealing meaningful relationships among ligand tokens that enabled us to enhance predictions for underrepresented ligands. We believe that our approach represents an essential step toward harnessing and upgrading large language models (LLMs) for robust and flexible protein prediction tasks.
@@ 3 Experiments
We evaluated Prot2Token in multiple tasks on different datasets, including the protein-level, residue-level, and protein-protein level. For a subset of these tasks, we incorporated a self-supervised pre-training stage for the autoregressive decoder as an initial step. In all experiments, the protein encoder in Prot2Token was initialized using the pre-trained ESM2-650m model. For the decoder part, we used an autoregressive language model with different configurations based on the size of the ESM encoder and hyperparameters of the autoregressive decoder (Appendix A.1). We only considered BARTSmiles as the chemical encoder for the protein-ligand affinity task and disabled it for the other tasks.
@@ 3.1 Classification
This category includes multi-class, multi-label and hierarchical classification tasks such as Deeploc 2.0 and ER. The results are shown in Tables 2 and 2. In Deeploc 2 dataset, we significantly improved the performance compared to the original method, and also, the ER task result showed that the performance boosted 7.5 percent by using multi-task learning. We could not calculate the Fmax metric for the EC and GO tasks, so we only considered the accuracy and F1 scores to evaluate performance. Consequently, direct comparisons with other methods were not possible. Supplementary results with additional details are in Appendix A.4.1.
@@
3.2 Regression
The results for these tasks are presented in Tables 4, 4, 6, and 6. Additional experimental details can be found in Appendix A.4.2. Across all regression tasks, Prot2Token consistently outperformed baseline methods from the PEER benchmark. Notably, in the fluorescence prediction task, multi-task learning led to a performance gain of up to 5.6% (Table 4). For mutation stability prediction, Prot2Token achieved a substantial improvement of over 51.5% compared to the best-performing baseline model as shown in Table 6.
@@ 3.2 Regression
3.3 Binding site
Method Spearman Encoder
PEER [27] (fine-tuned) 1.559 ESM1-1b
PEER [27] (fine-tuned) 1.562 ProtBert
@@ 3.3 Binding site
Table 6: Comparison of mutation effect prediction models on the ProteinGym benchmark with supervised 5-fold cross-validation. Additional baselines are included from the original ProteinGym paper [30].
Method Spearman
ESM-1v 0.542
MSAT 0.568
Tranception 0.571
ProteinNPT 0.613
Prot2Token-C 0.9294 $\pm$ 0.0018
Ligand 10 ligands $\dagger$ 10 ligands 20 ligands 30 ligands 41 ligands
Average 0.1883 0.6076 0.5942 0.6181 0.6132
Weighted Average 0.1849 0.6297 0.6277 0.6368 0.6353
We introduced a self-supervised pre-training phase for the decoder to enhance model initialization and improve predictive performance of binding site prediction-type tasks. This strategy significantly improves the model’s ability across different tasks, particularly benefiting those with a wide range of binding site indices. A detailed rationale and methodology for this self-supervised pre-training are provided in Appendix A.4.3. We report high-level performance results of protein-ligand binding site prediction in Table 7, demonstrating that Prot2Token achieves competitive predictive accuracy across various ligand types with the help of self-supervised pre-training (see detailed results of this task and protein-protein binding site in Appendix A.4.4).
Furthermore, to understand the representation learned by the task tokens, we explored their embeddings and identified relationships that correlate closely with biochemical properties (Appendix A.5). These findings, visualized in Figure 17 in the appendix, indicate that task tokens not only serve as input identifiers but also encode biologically relevant information. Leveraging these insights, we further utilized the learned relationships to boost predictive accuracy for underrepresented ligands, achieving significant performance gains as summarized in Table 34. More details in appendix A.5.5.
3.4 Sequence-to-sequence
Table 8: 3D structure prediction on continuous automated model evaluation (CAMEO) (Jan 2024 to Jan 2025) [35]. TM-score and inference time (one A100-80GB GPU) are reported for a representative 384-residue protein sequence. $\dagger$ Duo to computational cost, results for AF2 methods are reported from the ESM2 publication, using the CAMEO benchmark from April 2022 and June 2022.
Method TM-score A100 Wall-clock (384-aa) Speed-up
Prot2Token-D 0.54 1–2 s $\approx$1000$\times$
ESMFold (ESM2-3B) [36] 0.79 14.2 s 77$\times$
AF2 (w/o MSA) [21$\dagger$ 0.41 20–30 s 54$\times$
AF2 (w/ MSA) [21$\dagger$ 0.88 18–25 min 1$\times$
In this part, we evaluated Prot2Token on residue-wise sequences by formulating it as a sequence labeling task, where the model generates a discrete token for each residue in the input protein sequence. The main focus of this section is on the challenging task of sequence-to-3D structure prediction. Here, Prot2Token is trained to generate discrete 3D structure tokens from amino acid sequences using a vector quantized variational autoencoder (VQ-VAE) based representation for protein backbone coordinates. The results are summarized in Table 8, which reports TM-score and runtime for representative structure prediction methods. Notably, Prot2Token-D demonstrates a dramatic speed advantage, producing structure predictions for a typical 384-residue protein in 1–2 seconds on a single A100 GPU—approximately three orders of magnitude faster than AF2 with multiple sequence alignment (MSA) input, which typically requires 18–25 minutes for inference. This substantial speed-up makes Prot2Token particularly well-suited for large-scale or real-time structure generation scenarios. Representative examples of successful and unsuccessful 3D structure predictions are illustrated in Figure 4 and Figure 14, respectively. Additionally, as shown in Figure 15, the training process has not yet fully converged, indicating that further improvements in predictive performance are likely with extended training. The result of secondary structure is reported in Table 27.
@@ 4 Discussion
Our work introduces Prot2Token, a unified framework designed to address a broad spectrum of protein prediction tasks, ranging from protein-level and residue-level properties to complex protein-protein interactions. The core innovation lies in its versatile tokenization strategy, which demonstrates remarkable breadth in generalizing across these diverse predictive challenges by converting them into a unified next-token prediction problem. Prot2Token effectively utilizes the embedded capabilities of pre-trained encoder-style PLMs by attaching them to a general-purpose autoregressive predictor. This inherent generality facilitates robust multi-task learning, allowing a single model to concurrently learn and execute a variety of tasks, where the learning of certain tasks can have a synergistic effect (Tables 2, 4 and 4), boosting the performance of others, thereby streamlining the prediction pipeline and notably improving computational efficiency when deploying a single model to serve multiple predictive functions. This work also examines the current limitations of such systems (Appendix 5), particularly their weakness in tasks reliant on precise spatial information, such as binding site prediction, and introduces auxiliary self-supervised learning objectives (Appendix A.4.3) specifically for the decoder component of our architecture to address this challenge (Table 7). These auxiliary objectives are designed to instill specific inductive biases within the decoder, effectively teaching it to better capture the structural determinants crucial for accurate binding site predictions.
A significant implication of Prot2Token is its potential to distill the predictive power of numerous highly specialized models, including those for 3D structure prediction like AlphaFold2, into a single, efficient system. This consolidation could yield a substantial increase in prediction speed, potentially by three to four orders of magnitude, transforming current methodologies into high-throughput approaches (Table 8). This avenue of research is therefore anticipated to present numerous readily attainable and impactful opportunities, enabling rapid annotation and functional characterization of the ever-expanding protein sequence universe. Furthermore, the Task tokens central to our framework offer an avenue for interpretability; analyzing their embeddings can potentially uncover novel biochemical relationships between different protein properties or tasks (Figure 17). Separately, this work also contributes to the critical goal of making protein language models more intrinsically structure-aware (Table 28). In a broader sense, the tokenization scheme presented here represents a conceptual step towards empowering general-purpose autoregressive large language models with the capabilities required for sophisticated protein prediction tasks.
Looking ahead, we believe several future directions are worth exploring. A comprehensive multi-task training approach, encompassing an even wider array of biologically relevant tasks, will be crucial to fully realize the framework’s potential for generalization and synergistic learning. Investigating more sophisticated stochastic sampling methods, beyond simple greedy decoding, could unlock finer-grained control over predictions and potentially reveal a richer landscape of possible outcomes. Perhaps most compelling is the extension of this approach to protein design. The ability to seamlessly integrate protein prediction and generation within a single unified model opens the door to performing multiple in silico steps of the drug design and discovery pipeline, from target property prediction to novel candidate generation, all within one cohesive computational environment. This could dramatically accelerate the development and reduce the associated costs of new therapeutics and biomaterials.
5 Limitations
While Prot2Token demonstrates considerable versatility and strong performance across a range of protein prediction tasks, several limitations should be acknowledged. Firstly, the inherent nature of the autoregressive decoder initially presented challenges in tasks demanding precise spatial understanding, such as binding site prediction. Although the introduction of self-supervised decoder pre-training significantly mitigated this by instilling crucial inductive biases, the base architecture required this augmentation to achieve optimal performance on such spatially sensitive tasks.
Secondly, the aspiration of a truly universal model capable of concurrently learning an extremely broad spectrum of protein-related tasks is tempered by current computational realities. The experiments presented herein, while extensive, focused on combinations of a subset of tasks. Scaling to an exhaustive multi-task learning scenario encompassing all possible protein prediction types simultaneously would entail substantial computational resources, potentially limiting the practical scope of joint training in its most comprehensive form.
Furthermore, in sequence-to-sequence prediction tasks, we observed that the autoregressive decoder could occasionally produce output sequences with lengths inconsistent with the input protein sequence (1 to 3 amino acids). This necessitated the implementation of post-inference constraints on the end-of-sequence token to ensure coherent and correctly dimensioned outputs, indicating that raw decoder outputs may not always perfectly align with expected structural constraints without such measures.
The performance of Prot2Token is also intrinsically linked to the capabilities of the foundational pre-trained protein encoders it utilizes for generating input embeddings. Any limitations or biases present in these underlying encoder models could invariably propagate and influence the predictive accuracy of the Prot2Token framework. Additionally, like many machine learning models, Prot2Token’s performance can be challenged by datasets exhibiting severe class imbalances. While the framework showed robust results on diverse benchmarks, tasks with highly skewed label distributions may require further specialized strategies to ensure equitable learning and prediction across all classes.
Finally, our current exploration of decoding strategies has primarily focused on greedy decoding. A more comprehensive investigation into various stochastic sampling methods (e.g., top-k, nucleus sampling) could potentially unlock finer-grained control over the generation process and reveal a richer landscape of possible predictions, but this remains an avenue for future work.

### v2

@@ Abstract
[…]
@@ 1 Introduction
While PLMs have significantly advanced protein-prediction tasks, current models require task-specific specialization after pre-training [5, 6]. This reliance on separate modules for distinct tasks leads to inefficient computational resource use and limited scalability. Most PLMs undergo post-training alignment with specialized predictor architectures for individual tasks, requiring independent training and fine-tuning—a time-consuming and resource-intensive approach [7]. A unified tokenization protocol capable of standardizing diverse protein-prediction tasks would overcome this limitation, streamlining protein function prediction and enhancing its accessibility for real-world applications.
To the best of our knowledge, despite the emergence of foundation models for proteins, no comprehensive strategy exists to systematically map these heterogeneous outputs into a shared generative space. Instead, researchers often modify existing foundation models to suit particular applications [8], such as predicting 3D protein structures from sequences using customized techniques [9, 10]. One key limitation is that most existing models are based on BERT-style architectures [11], while effective for providing meaningful representation, lack the flexibility needed for diverse and controllable prediction capabilities. In natural language processing (NLP), the transition from BERT-style models to autoregressive GPT-style models has enabled more dynamic and human-understandable instructions (prompts) to control the generation process and therefore, handling a diverse set of predictions within the NLP domain [12, 13]. A similar paradigm shift is necessary in protein research, moving beyond static encoders toward more advanced generative AI approaches that provide more comprehensive predictive capabilities.
@@
To address these limitations, this work takes a significant step toward unifying protein prediction by establishing a universal tokenization protocol that categorizes diverse tasks into five sets. We introduce a universal protocol for tokenizing different protein-prediction tasks, enabling a general autoregressive transformer predictor to leverage existing BERT-style PLMs (Figure 1). This generative approach, guided by a unified next-token prediction loss, demonstrates generality across multiple protein-prediction task categories, including protein-level, residue-level, and protein-protein interaction-level tasks. We illustrate its versatility through extensive evaluation on five categories of tasks: Classification, Regression, Binding Site, Sequence-to-Sequence, and Other. Specific examples evaluated include kinase phosphorylation site prediction, protein-ligand binding site prediction, protein 3D structure prediction, and protein mutation stability assessment. Furthermore, our framework inherently supports multi-task learning, and we provide initial analyses demonstrating synergistic performance improvements when related tasks are trained jointly.
For certain specialized tasks, such as predicting binding sites, we show that initializing the decoder through self-supervised pre-training significantly boosts performance. Specifically, for protein-ligand binding site prediction, we further analyzed the learned token representations, revealing meaningful relationships among ligand tokens that enabled us to enhance predictions for underrepresented ligands. We believe that our approach represents an essential step toward harnessing and upgrading large language models (LLMs) for robust and flexible protein prediction tasks.
@@ 3 Experiments
We evaluated Prot2Token in multiple tasks on different datasets, including the protein-level, residue-level, and protein-protein level. For a subset of these tasks, we incorporated a self-supervised pre-training stage for the autoregressive decoder as an initial step. In all experiments, the protein encoder in Prot2Token was initialized using the pre-trained ESM2-650m model. For the decoder part, we used an autoregressive language model with different configurations based on the size of the ESM encoder and hyperparameters of the autoregressive decoder (Appendix A.1). We only considered BARTSmiles as the chemical encoder for the protein-ligand affinity task and disabled it for the other tasks. We optimized the number of unfrozen ESM-2 encoder layers for each task to align model capacity with task complexity and data availability; specific hyperparameters for each task are detailed in Appendix A.4.
@@ 3.1 Classification
This category includes multi-class, multi-label and hierarchical classification tasks such as Deeploc 2.0 and ER. The results are shown in Tables 2 and 2. In Deeploc 2 dataset, we significantly improved the performance compared to the original method, and also, the ER task result showed that the performance was boosted 7.5 percent by using multi-task learning. We could not calculate the Fmax metric for the EC and GO tasks, so we only considered the accuracy and F1 scores to evaluate performance. Consequently, direct comparisons with other methods were not possible. Supplementary results with additional details are in Appendix A.4.1.
3.2 Regression
@@
The results for these tasks are presented in Tables 4, 4, 7, and 7. Additional experimental details can be found in Appendix A.4.2. Across all regression tasks, Prot2Token consistently outperformed baseline methods from the PEER benchmark. Notably, in the fluorescence prediction task, multi-task learning led to a performance gain of up to 5.6% (Table 7). For mutation stability prediction, Prot2Token achieved a substantial improvement of over 51.5% compared to the best-performing baseline model as shown in Table 4.
3.3 Binding site
Method RMSE Encoder
PEER [27] (fine-tuned) 1.559 ESM1-1b
PEER [27] (fine-tuned) 1.562 ProtBert
[…]
Table 4: Comparison of mutation effect prediction models on the ProteinGym benchmark with original supervised 5-fold cross-validation indices. Additional baselines are included from the original ProteinGym paper [30]. $\dagger$ denotes a linear layer fine-tuned on the last four encoder blocks.
Method Spearman
ESM‑1v 0.542
MSAT 0.568
Tranception 0.571
ProteinNPT 0.613
[…]
Prot2Token-C 0.9294 $\pm$ 0.0018
Ligand 10 ligands $\dagger$ 10 ligands 20 ligands 30 ligands 41 ligands
Average 0.1883 0.6076 0.5942 0.6181 0.6132
Weighted Average 0.1849 0.6297 0.6277 0.6368 0.6353
We introduced a separate self-supervised pre-training stage for the decoder weights to enhance model initialization to improve predictive performance of binding site prediction-type tasks before training such tasks. This strategy significantly improves the model’s ability across tasks require a wide range of binding site indices. A detailed rationale and methodology for this self-supervised pre-training are provided in Appendix A.4.3. We reported high-level performance results of protein-ligand binding site prediction in Table 5, demonstrating that Prot2Token achieves competitive predictive accuracy across various ligand types with the help of self-supervised pre-training (see detailed results of this task and protein-protein binding site in Appendix A.4.4).
@@ 3.3 Binding site
Furthermore, to understand the representation learned by the task tokens, we explored their embeddings and identified relationships that correlate closely with biochemical properties (Appendix A.5). These findings, visualized in Appendix, Figure 17, indicate that task tokens not only serve as input identifiers but also encode biologically relevant information. Leveraging these insights, we further utilized the learned relationships to boost predictive accuracy for underrepresented ligands, achieving significant performance gains as summarized in Table 34. More details in Appendix A.5.5.
3.4 Sequence-to-sequence
In this part, we evaluated Prot2Token on residue-wise sequences by formulating it as a sequence labeling task, where the model generates a discrete token for each residue in the input protein sequence. The main focus of this section is on the challenging task of sequence-to-3D structure prediction. Here, Prot2Token is trained to generate discrete 3D structure tokens from amino acid sequences using a vector quantized variational autoencoder (VQ-VAE) based representation for protein backbone coordinates. The results are summarized in Table 8, which reports TM-score and runtime for representative structure prediction methods. Notably, Prot2Token-D demonstrates a dramatic speed advantage, producing structure predictions for a typical 384-residue protein in 1–2 seconds on a single A100 GPU—approximately three orders of magnitude faster than AF2 with multiple sequence alignment (MSA) input, which typically requires 18–25 minutes for inference. This substantial speed-up makes Prot2Token particularly well-suited for large-scale or real-time structure generation scenarios. Representative examples of successful and unsuccessful 3D structure predictions are illustrated in Figure 13 and Figure 14, respectively. Interestingly, although the validation perplexity continues to decrease (Figure 15), structure accuracy plateaus at $\sim$0.55 TM-score on CAMEO 2024; this aligns with the $\sim$0.60 TM-score reconstruction ceiling of the VQ-VAE tokenizer, indicating a tokenizer-imposed bottleneck rather than a lack of decoder convergence. Results for secondary structure appear in Appendix; Table 27.
Table 8: 3D structure prediction on continuous automated model evaluation (CAMEO 2024) (Jan 2024 to Jan 2025) [35]. Inference time of all methods is reported on identical A100 hardware for a representative 384-residue protein sequence. $\dagger$ Due to computational cost, TM-score for AF2 methods is reported from the ESM2 publication, using the CAMEO benchmark from April 2022 and June 2022.
Method TM‑score A100 Wall‑clock (384‑aa) Speed‑up
@@ 3.4 Sequence-to-sequence
[…]
@@ 4 Discussion
This work introduces Prot2Token, a unified tokenization framework that reimagines protein prediction tasks as a next token prediction. By developing a versatile tokenization strategy, we demonstrate that a single autoregressive decoder can effectively map the latent representations of pre-trained PLMs to a diverse array of biological outputs—ranging from residue-level annotations and scalar properties to complex 3D structural coordinates. This approach has the potential to represent a paradigm shift from utilizing specialized, task-specific heads to employing a general-purpose sequence generation mechanism, thereby enabling multi-task learning across completely different protein tasks that seemed impossible before.
4.1 Key Insights and Empirical Observations
Throughout the development and evaluation of Prot2Token, we observed several distinct behaviors that highlight both the capabilities and the idiosyncrasies of modeling proteins via next-token prediction.
Unified Tokenization. A central outcome of this study is the validation of our universal tokenization protocol. Our primary objective was not to train a single, monolithic model covering all possible downstream tasks but rather to demonstrate the feasibility of mapping the vast landscape of protein prediction problems into a cohesive generative framework. As illustrated in Figure 3, we established that virtually all protein tasks can be categorized into five structural categories: Classification, Regression, Binding Site, Sequence-to-Sequence, and Other (complex composite) tasks. By selecting and evaluating at least one representative candidate from each category, we confirmed that this unified next-token prediction format yields robust performance comparable to, and often exceeding, specialized or baseline models. This provides a scalable blueprint for bringing diverse biological tasks into the generative interface without requiring bespoke architectures for each domain.
Hierarchical Regression. A distinct advantage emerged from our single-digit tokenization strategy for regression tasks, which benefited most significantly from the unified framework. Unlike standard predictors that output a continuous value in a single "shot," our autoregressive approach effectively performs a coarse-to-fine prediction. By generating values digit-by-digit, the model first establishes the order of magnitude before iteratively refining the precision. This multi-step process allows for dynamic internal adjustments as the prediction becomes more granular. Consequently, Prot2Token outperformed the strong ESM + linear probe baseline—utilizing the same encoder—in nearly all regression benchmarks. As evidenced by Tables 4, 4, 7, 7, and Appendix Tables 21, 22, our method significantly surpasses the best baseline in all tasks except protein thermostability prediction (Table 23), validating that discretizing continuous spaces into hierarchical token sequences is a highly effective modeling strategy.
Multi-Task Learning. While investigating the full spectrum of cross-task learning (including negative transfer) was constrained by the computational cost of balancing highly diverse data distributions, we utilized protein-ligand binding site prediction as a controlled environment to study these dynamics. By treating the prediction of binding sites for 41 distinct ligands as separate tasks (each defined by a unique prompt) we observed that aggregating tasks yielded clear synergistic benefits, particularly for data-scarce targets. Crucially, our deep investigation into the learned task token embeddings (Appendix A.5) revealed that the model encoded physicochemical correlations, grouping ligands by properties such as molecular weight and charge without explicit supervision. This semantic structure in the task space facilitated knowledge transfer, allowing the model to leverage latent patterns from over-represented ligands to significantly boost prediction accuracy for chemically similar, underrepresented ligands (Table 34). Beyond the ligand domain, we similarly observed beneficial correlations when aggregating distinct prediction types to form auxiliary tasks for a specific target, as demonstrated by the performance gains reported in Tables 2 and 7.
[…]
4.2 Limitations
The quality and distribution of labels vary significantly across protein prediction tasks. While some datasets are uniform, others suffer from extreme imbalance; for instance, the Fold classification dataset contains classes with single samples (Appendix A.4.1), and binding site indices follow a severe long-tail distribution (Figure 9 and Figure 10). Our analysis suggests this sensitivity is a data limitation rather than an architectural flaw, as Prot2Token excels when data is abundant. We mitigated some of these irregularities via engineering interventions—such as token weighting and self-supervised pre-training—rather than fundamental architectural changes. This heterogeneity necessitated validating the tokenization protocol across task categories rather than pursuing a monolithic “all-task” training run. However, we anticipate this issue will diminish in real-world applications where datasets are typically magnitudes larger than academic benchmarks, thereby reducing the prevalence of extreme data label sparsity.
Furthermore, performance is intrinsically bounded by the foundational models used. Biases in the protein encoder (e.g., ESM2) propagate to predictions, for instance, the accuracy plateau in 3D structure prediction (TM-score $\approx$ 0.55) reflects the reconstruction ceiling of the VQ-VAE tokenizer. Consequently, closing these gaps requires integrating higher-fidelity components rather than altering the predictor architecture.
4.3 Future Directions
Looking ahead, several research avenues promise to extend the capabilities of this framework. A primary technical objective is the development of high-fidelity, discrete tokenizers for 3D structures that can surpass the current reconstruction bottlenecks, potentially allowing the speed of Prot2Token to be paired with high-accuracy folding. Additionally, moving beyond deterministic greedy decoding to explore stochastic sampling strategies could unlock a richer landscape of probabilistic outputs, which is particularly valuable for modeling conformational diversity in structure prediction.
Perhaps the most compelling direction is the inversion of the current paradigm: extending Prot2Token from prediction to generation. The unified architecture naturally supports both workflows where the model could not only predict properties from sequence but also generate novel sequences conditioned on desired property tokens. This would allow for the seamless integration of prediction and design within a single cohesive model, potentially accelerating the in silico development of novel therapeutics and biomaterials.